# LoRA is All You Need for Safety Alignment of Reasoning LLMs

### v3

@@ Abstract
Reasoning LLMs have demonstrated remarkable breakthroughs in solving complex problems that were previously out of reach. To ensure LLMs do not assist with harmful requests, safety alignment fine-tuning is necessary in the post-training phase. However, safety alignment fine-tuning has recently been shown to significantly degrade reasoning abilities, a phenomenon known as the “Safety Tax”. In this work, we show that using LoRA for SFT on refusal datasets effectively aligns the model for safety without harming its reasoning capabilities. This is because restricting the safety weight updates to a low-rank space minimizes the interference with the reasoning weights. Our extensive experiments across four benchmarks covering math, science, and coding show that this approach produces highly safe LLMs—with safety levels comparable to full-model fine-tuning—without compromising their reasoning abilities. Our ablation studies further identify three key factors in LoRA: (1) rank-$1$ updates are sufficient to achieve the best reasoning and safety performance, (2) the up projection layers are the most critical modules, with LoRA applied to them alone achieving even better results, and (3) middle layers are more effective than early or late layers. Together, these findings show that strong safety and reasoning can be achieved at minimal computational cost when updates are applied in the right places. Additionally, we observe that LoRA induces weight updates with smaller overlap with the initial weights compared to full-model fine-tuning. Finally, while our attempts to further reduce this overlap yield only modest improvements on some tasks, they highlight the potential of developing methods that more reliably optimize the reasoning–safety tradeoff. Our source code is available at https://github.com/YihaoXue/lora-safety-reasoning.
@@ 1 Introduction
Large language models (LLMs) have made remarkable progress across a wide range of tasks. A major recent breakthrough is the emergence of LLMs with advanced reasoning capabilities, enabling them to solve complex problems previously out of reach. However, recent studies have reported significant safety risks associated with reasoning-capable models (Jiang et al., 2025; Zhou et al., 2025; Huang et al., 2025; Li et al., 2025a). Indeed, reasoning fine-tuning—the process through which LLMs acquire these capabilities—often compromises safety, even when starting from a safety-aligned checkpoint (Jiang et al., 2025; Zhou et al., 2025; Zhao et al., 2025; Li et al., 2025a). For example, Jiang et al. (2025) show that models distilled for reasoning from DeepSeek-R1 become substantially less safe than their original base models.
@@
There has been significant effort in the literature to preserve LLMs’ safety alignment during instruction fine-tuning. However, these approaches are not applicable to reasoning fine-tuning. First, reasoning fine-tuning datasets are often highly curated (Muennighoff et al., 2025) and unlikely to contain unsafe content. Thus, data filtering methods such as those proposed by Shen et al. (2024); Choi et al. (2024); Bianchi et al. (2023) are not applicable. In addition, methods that restrict model updates during fine-tuning (Hsu et al., 2024; Mukhoti et al., 2023) are ineffective in the reasoning setting, as acquiring reasoning capabilities typically requires longer training and more substantial weight updates compared to instruction fine-tuning. To the best of our knowledge, the current literature does not offer any method for safety alignment of reasoning models.
Instead, the prevailing strategy is to apply a secondary safety alignment phase after reasoning capabilities have been acquired. This phase—often implemented via supervised fine-tuning (SFT) or reinforcement learning (RL)—has become a standard step in modern LLM development. Although safety alignment fine-tuning can significantly improve model safety, it often comes at a steep cost to reasoning performance—a phenomenon referred to as the “Safety Tax” (Huang et al., 2025). Even incorporating chain-of-thought (CoT) style reasoning into safety fine-tuning datasets (Jiang et al., 2025) cannot succeed in fully preserving reasoning abilities (Huang et al., 2025).
@@
In this work, we investigate the algorithmic factors that contribute to this trade-off. Existing evidence suggests that safety-related behavior in LLMs is often governed by a small number of dominant directions—either in activation space, such as steering vectors (Panickssery et al., 2023) or refusal features (Arditi et al., 2024; Yu et al., 2024), or in weight space. In particular, Jain et al. (2024); Wei et al. (2024) show that safety-critical weights tend to lie in a low-rank subspace. In our analysis, we find that the model undergoes relatively high-rank changes during full-model fine-tuning (see Figure 1), which results in Safety Tax. This highlights a key insight: although achieving safety may require modifying weights only along a low-rank subspace, full-model fine-tuning permits arbitrary updates, potentially introducing many unnecessary changes that interfere with reasoning.
@@
Our extensive experiments reveal the surprising effectiveness of a simple recipe for safety alignment of reasoning models: Applying LoRA during SFT using a straightforward direct refusal data set. Despite its simplicity, this approach achieves safety performance on par with full-model alignment, while preserving reasoning capability close to that of the original reasoning-tuned model. This result holds for both 7B and 14B models and is validated across four benchmarks spanning mathematics, science, and code generation. It represents a rare “one stone, three birds” outcome: strong safety, strong reasoning, and computational efficiency.
@@
Moreover, we further ablate the LoRA configuration to understand “how much LoRA is sufficient”. We make three key findings. (1) Setting the rank to $r=1$ achieves the best reasoning–safety tradeoff (in terms of the Pareto frontier when the number of training epochs is varied). This is encouraging, as it shows that strong performance on both reasoning and safety can be achieved at the lowest possible fine-tuning cost. (2) Updating only the up projection layers in the MLP yields an even better tradeoff than updating the full MLP, while updating only the gate or down projections performs worse. This highlights the central role of the up projection and motivates future research into understanding why it is so effective. (3) Middle layers are more important for a good reasoning–safety tradeoff: updating only 16 middle layers is often sufficient, whereas updating early or late layers yields worse results. Interestingly, this connects to prior findings that safety-critical features often emerge in the middle layers of LLMs (Panickssery et al., 2023; Arditi et al., 2024). Overall, our results provide meaningful insights into the key elements of LoRA configuration that affect the reasoning–safety tradeoff, and can help achieve strong performance at minimal cost.
Additionally, we explore the weight structure imposed by LoRA to understand the differences it introduces. We find that LoRA updates are not only low-rank by design but also exhibit smaller alignment with the weights of the original reasoning model compared to those from full-model fine-tuning—across most layers. While the reduction in alignment is small, it may suggest that LoRA updates are less disruptive to reasoning-related weights. We further investigate whether explicitly reducing such overlap—via regularization or post-hoc weight merging—can improve safety or reasoning capabilities. We find that one post-hoc method for reducing overlap achieves a modest gain in the reasoning–safety trade-off on some tasks. This shows promise, but more effort is needed to develop approaches that yield consistent improvements across tasks, which we consider a valuable direction for future work.
@@ 2 Related Work
To develop LLMs that are both safe and capable, models can be safety aligned before or after fine-tuning.
@@
Fine-tuning a safety-aligned model. Fine-tuning a safety-aligned model often leads to safety degradation. For instruction fine-tuning, safety degradation is shown across various model architectures and optimization strategies, including full-model and LoRA fine-tuning (Qi et al., 2023; Hsiung et al., 2025; Zhan et al., 2023). To mitigate this issue, several methods have been proposed. Shen et al. (2024); Choi et al. (2024) focus on data filtering, aiming to remove unsafe examples from the fine-tuning data. Bianchi et al. (2023) shows that injecting just a few hundred safety examples during instruction fine-tuning can improve safety. Peng et al. (2025) leverages an existing guardrail model to encourage safe response segments while suppressing unsafe ones. Lyu et al. (2024) emphasizes the importance of prompt templates in preserving safety. Algorithmic approaches have also been explored: Hsu et al. (2024) propose projecting LoRA updates into a “safety subspace” derived from differences between aligned and unaligned models, while Mukhoti et al. (2023) introduce regularization techniques to constrain changes in intermediate representations during fine-tuning.
This approach is, however, not applicable to reasoning models. This is because acquiring reasoning capabilities typically requires longer training and more substantial weight updates compared to instruction fine-tuning. This results in losing the initial safety alignment of the model. Indeed, several recent studies have reported significant safety risks associated with reasoning-capable models Jiang et al. (2025); Zhou et al. (2025); Huang et al. (2025); Li et al. (2025a).
Safety alignment after fine-tuning. Aligning a fine-tuned model is typically done using supervised fine-tuning (SFT) and/or reinforcement learning (RL) (Wei et al., 2021; Griffith et al., 2013; Dai et al., 2023; Ouyang et al., 2022; Rafailov et al., 2023; Bai et al., 2022; Guan et al., 2024). However, this approach introduces a key trade-off: safety alignment can substantially impair model performance. Huang et al. (2025) characterizes this as the “Safety Tax,” showing that safety-aligned models often lose much of their reasoning ability. Jiang et al. (2025) attempted to address this issue by constructing a safety fine-tuning dataset with long chain-of-thought (CoT) responses, but the resulting models still showed noticeable drops in reasoning performance (Huang et al., 2025).
In our work, we show that a simple application of LoRA for safety alignment of reasoning models can effectively align reasoning-capable models without compromising their performance.
@@ 3 LoRA for Safety Alignment Without Compromising Reasoning
In this section, we investigate whether the “Safety Tax” (Huang et al., 2025) can be mitigated. In particular, we aim to answer the following question: can we align a model for safety without compromising its reasoning capabilities? We focus on the same setup as in (Huang et al., 2025), where SFT is performed on safety datasets that provide harmful requests paired with refusal responses.
Our key observation is that during full-model fine-tuning, which is used in (Huang et al., 2025), the weights undergo relatively high-rank changes. As shown in Figure 1, we observe high stable ranks for the weight updates, i.e., differences between the fine-tuned model’s weights and those of the initial model, in most layers. However, prior evidence suggests that safety behavior in LLMs is typically governed by only a single or a few directions in the activations (Panickssery et al., 2023; Arditi et al., 2024) and weights (Wei et al., 2024; Jain et al., 2024), indicating that a small low-rank modification may be sufficient to induce safe behavior, without altering the entire weight space. Thus, we conjecture that the degradation in reasoning performance is caused by full-model fine-tuning introducing unnecessary changes in many directions, which interfere with critical weights responsible for reasoning.
To address this, we consider Low-Rank Adaptation (LoRA) (Hu et al., 2022), originally proposed as a parameter-efficient fine-tuning method to reduce training cost and memory usage. Rather than updating the full weight matrices, LoRA injects trainable low-rank matrices into existing layers while keeping the original weights frozen. Formally, a weight matrix ${\bm{W}}\in\mathbb{R}^{d\times k}$ is modified as:
@@
LoRA is particularly well-suited to our needs: it restricts updates to a low-rank subspace, thereby significantly reducing interference with the original weights. We will show in our experiments that this method works excellently, enabling the model to become safe while maintaining strong reasoning performance across benchmarks. As an additional benefit, LoRA is significantly more computationally efficient than full-model fine-tuning.
@@ 4 LoRA Bypasses the “Safety Tax”
In this section, we first introduce our safety fine-tuning and evaluation pipeline. Then, we evaluate models’ safety alignment and reasoning performance after full-model and LoRA safety fine-tuning.
@@ 4.1 Safety Alignment Fine-Tuning of Reasoning LLMs
We begin with a reasoning-capable language model and perform safety alignment fine-tuning. Specifically, we apply supervised fine-tuning (SFT) on a safety dataset consisting of harmful questions paired with refusal responses, aiming to teach the model to reject harmful requests. We choose SFT over reinforcement learning (RL)-based techniques because it is simpler, less expensive, and does not require additional components such as a reward model. However, we expect our results to apply to RL safety alignment as well.
In our training setup, we compare two approaches: (1) full-model fine-tuning, as in (Huang et al., 2025), where all model parameters are updated using standard gradient-based optimization; and (2) LoRA fine-tuning, as described in Section 3.
4.2 Evaluation of the Fine-Tuned Model
After safety alignment fine-tuning is completed, we evaluate two aspects of the model: (1) safety, which is assessed using a dataset of harmful questions. We sample responses from the model for these questions and use Llama-Guard-3-8B—a model specialized in safety evaluation and shown to be the strongest safety evaluator in (Jiang et al., 2025)—to determine whether the responses are safe. The safety score is defined as the proportion of questions for which the model’s response is judged to be harmful. (2) reasoning ability, evaluated using multiple standard benchmark datasets containing questions on math, science, and coding—widely used to assess models’ reasoning capabilities. We consider the commonly used metric Pass@1 to measure accuracy on these benchmarks. For each question, we sample $n$ responses, compute the fraction of correct responses, and then average this accuracy over all questions. We set $n=8$.

### v4

@@ Abstract
Reasoning-capable LLMs have achieved major breakthroughs in solving complex problems, but recent work shows that acquiring and deploying strong reasoning can introduce significant safety risks. A common mitigation is to apply a secondary safety-alignment phase after reasoning is learned; however, safety alignment often degrades reasoning performance—a phenomenon known as the “Safety Tax”. In this work, we show that a simple approach can largely bypass this trade-off: applying LoRA during SFT on refusal datasets. Despite its simplicity, this recipe achieves safety comparable to full-model alignment while preserving reasoning performance close to the original reasoning-tuned model, and the result holds across multiple model sizes and architectures, two safety benchmarks, and four reasoning benchmarks spanning mathematics, science, and code generation. We further ablate LoRA configurations and find that (1) rank-$1$ updates are sufficient to achieve the best safety–reasoning trade-off, (2) applying LoRA only to the MLP up-projection layers can outperform updating the full MLP, and (3) updating middle layers is more effective than updating early or late layers. Finally, we provide a theoretical analysis that helps understand when and why LoRA works, revealing that overshooting the rank budget (using a larger rank than needed for the finetuning task) induces base-task degradation at a rate inversely proportional to the intrinsic dimensionality of the base task. This suggests LoRA is most effective when the finetuning task is low-rank and the base capability is high-rank.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have made remarkable progress across a wide range of tasks. A major recent breakthrough is the emergence of LLMs with advanced reasoning capabilities, enabling them to solve complex problems previously out of reach. However, recent studies have reported significant safety risks associated with reasoning models (Jiang et al., 2025; Zhou et al., 2025; Huang et al., 2025; Li et al., 2025a). Indeed, reasoning finetuning—the process through which LLMs acquire these capabilities—often compromises safety, even when starting from a safety-aligned checkpoint (Jiang et al., 2025; Zhou et al., 2025; Zhao et al., 2025; Li et al., 2025a). For example, Jiang et al. (2025) show that models distilled for reasoning from DeepSeek-R1 become substantially less safe than their original base models.
@@
There has been significant effort in the literature to preserve LLMs’ safety alignment during instruction finetuning. However, these approaches are not directly applicable to reasoning finetuning. First, reasoning finetuning datasets are often highly curated (Muennighoff et al., 2025) and unlikely to contain unsafe content, limiting the applicability of data filtering methods (Shen et al., 2024; Choi et al., 2024; Bianchi et al., 2023). Second, methods that restrict model updates during finetuning (Hsu et al., 2024; Mukhoti et al., 2023) are often ineffective in the reasoning setting, where acquiring reasoning capabilities typically requires longer training and more substantial weight updates than instruction finetuning. To the best of our knowledge, the current literature does not offer a practical method for safety-preserving reasoning finetuning. Instead, the prevailing strategy is to apply a secondary safety alignment phase after reasoning capabilities have been acquired. This phase—often implemented via supervised finetuning (SFT) or reinforcement learning (RL)—has become a standard step in modern LLM development. Although safety alignment finetuning can significantly improve model safety, it often comes at a steep cost to reasoning performance, a phenomenon referred to as the “Safety Tax” (Huang et al., 2025). Even incorporating chain-of-thought (CoT) style reasoning into safety finetuning datasets (Jiang et al., 2025) cannot fully preserve reasoning abilities (Huang et al., 2025).
@@
In this work, we investigate the algorithmic and structural factors underlying this trade-off. Existing evidence suggests that safety-related behavior in LLMs is often governed by a small number of dominant directions—either in activation space (e.g., steering vectors (Panickssery et al., 2023) and refusal features (Arditi et al., 2024; Yu et al., 2024)) or in weight space (e.g., Jain et al. (2024); Wei et al. (2024) show that safety-critical weights tend to lie in a low-rank subspace). At the same time, we find that full-model safety alignment induces relatively high-rank changes across layers (Figure 1), which can unnecessarily perturb the reasoning solution and contribute to Safety Tax. This motivates a simple hypothesis: if safety can be achieved by modifying weights along a low-dimensional subspace, then restricting updates to that subspace should reduce interference with reasoning.
@@
A simple recipe: LoRA for safety alignment of reasoning models. Guided by this hypothesis, we find the surprising effectiveness of a simple recipe for safety alignment of reasoning models: applying LoRA during SFT using a straightforward direct-refusal dataset. Despite its simplicity, this approach achieves safety performance on par with full-model alignment while preserving reasoning capability close to that of the original reasoning-tuned model. This result holds across three model sizes (7B, 8B, and 14B), two architectures (Qwen and Llama), two safety benchmarks covering a variety of harm categories, and four reasoning benchmarks spanning mathematics, science, and code generation. It represents a rare “one stone, three birds” outcome: strong safety, strong reasoning, and computational efficiency.
@@
Understanding how much LoRA is sufficient. We further ablate the LoRA configuration to understand “how much LoRA is sufficient”. We make three findings. (1) Setting the rank to 1 is sufficient (in terms of the Pareto frontier as the number of training epochs varies), so strong trade-off can be achieved at the lowest possible finetuning cost. (2) Updating only the up-projection layers in the MLP yields an even better trade-off than updating the full MLP, while updating only the gate or down projections performs worse. (3) Middle layers are more important for a good reasoning–safety trade-off: updating only 16 middle layers is often sufficient, whereas updating early or late layers yields worse results. Interestingly, this connects to prior findings that harmfulness-related representations often emerge in the middle layers of LLMs (Panickssery et al., 2023; Arditi et al., 2024). Overall, these results provide practical guidance on LoRA configurations that achieve strong performance at minimal cost.
Theory: when and why can LoRA avoid Safety Tax? Finally, we provide a theoretical explanation for when LoRA should be expected to succeed. We analyze a stylized linear regression model with a base task (capturing the capability to preserve) and a finetuning task (capturing safety), and compare full-model finetuning to LoRA. The theory captures the phenomenon that full-model finetuning fits the finetuning task but forgets the base task, whereas LoRA can perform well on both tasks. Moreover, it yields two key insights about LoRA. First, LoRA can fit the finetuning task well only when the dimensionality of the finetuning objective is comparable to the rank budget $r$. Second, when $r$ exceeds the effective dimension required to represent the finetuning task, the resulting overshoot (i.e., choosing $r$ larger than necessary) induces base-task degradation that scales roughly as
$\begin{aligned} &\text{base-task}\\[-3.01385pt] &\text{degradation}\end{aligned}\propto\frac{\text{overshoot}}{\text{intrinsic dimension of the base capability}}.$
In other words, higher-dimensional base capabilities are substantially more tolerant to overshoot: increasing $r$ beyond what is required has a much smaller impact on base performance when the base capability is intrinsically high-dimensional. Together, these results predict that LoRA is most effective when the finetuning objective is approximately low-dimensional while the base capability is intrinsically high-dimensional, making it tolerant to the inevitable mismatch in choosing $r$ in practice. We validate these predictions both through controlled numerical experiments and in a realistic setting. When safety-finetuning an instruction-tuned model (where the base capability is instruction following rather than reasoning), LoRA is no longer effective, consistent with the theory’s requirement that the preserved base capability be intrinsically high-dimensional.
@@ 2 Related Work
To develop LLMs that are both safe and capable, models can be safety aligned before or after finetuning.
@@
Finetuning a safety-aligned model. Finetuning a safety-aligned model often leads to safety degradation. For instruction finetuning, safety degradation is shown across various model architectures and optimization strategies, including full-model and LoRA finetuning (Qi et al., 2023; Hsiung et al., 2025; Zhan et al., 2023). To mitigate this issue, several methods have been proposed. Shen et al. (2024); Choi et al. (2024) focus on data filtering, aiming to remove unsafe examples from the finetuning data. Bianchi et al. (2023) shows that injecting just a few hundred safety examples during instruction finetuning can improve safety. Peng et al. (2025) leverages an existing guardrail model to encourage safe response segments while suppressing unsafe ones. Lyu et al. (2024) emphasizes the importance of prompt templates in preserving safety. Algorithmic approaches have also been explored: Hsu et al. (2024) propose projecting LoRA updates into a “safety subspace” derived from differences between aligned and unaligned models, while Mukhoti et al. (2023) introduce regularization techniques to constrain changes in intermediate representations during finetuning. These approaches are, however, not applicable to reasoning finetuning. Acquiring reasoning capabilities typically requires longer training and more substantial weight updates compared to instruction finetuning. This results in losing the initial safety alignment of the model. Indeed, several recent studies have reported significant safety risks associated with reasoning-capable models (Jiang et al., 2025; Zhou et al., 2025; Huang et al., 2025; Li et al., 2025a).
Safety alignment after finetuning. Aligning a fine-tuned model is typically done via supervised finetuning (SFT) and/or reinforcement learning (RL) (Wei et al., 2021; Griffith et al., 2013; Dai et al., 2023; Ouyang et al., 2022; Rafailov et al., 2023; Bai et al., 2022; Guan et al., 2024). For reasoning models, Huang et al. (2025) identifies a key trade-off—termed the “safety tax”—where safety alignment can substantially degrade reasoning ability. Jiang et al. (2025) construct a safety finetuning dataset with long CoT responses, but the resulting models still show noticeable drops in reasoning performance (Huang et al., 2025). Related work also studies how to mitigate over-refusal in safety-aligned models, such as post-hoc model editing (Dabas et al., 2025) or inference-time methods (Ray and Bhalani, 2024; Shi et al., 2024; Wang et al., 2024a). However, these approaches do not directly address the trade-off induced by safety finetuning itself, and they primarily focus on instruction-following behavior rather than reasoning-specific capabilities.
In our work, we show that a simple application of LoRA can effectively align reasoning models for safety without compromising their reasoning performance. We note that our work is specific to safety alignment of reasoning models, and should not be confused with either utility finetuning of an already safety-aligned model or safety alignment of non-reasoning models. The surprising effectiveness of LoRA appears to be tied to this specific setup and does not manifest as strongly in other settings, due to the interaction between LoRA and the tasks’ inherent dimensionalities, as analyzed in our theory (Sec. 6.1) and empirically observed in Fig. 9.
@@ 3 LoRA for Safety Alignment Without Compromising Reasoning
We investigate whether the “safety tax” (Huang et al., 2025) can be mitigated—i.e., whether a model can be safety-aligned without sacrificing reasoning ability. We follow the setup of Huang et al. (2025), performing SFT on safety datasets that pair harmful requests with refusal responses.
Our key observation is that during full-model finetuning, which is used in (Huang et al., 2025), the weights undergo relatively high-rank changes. As shown in Figure 1, we observe high stable ranks for the weight updates, i.e., differences between the fine-tuned model’s weights and those of the initial model, in most layers. However, prior evidence suggests that safety behavior in LLMs is typically governed by only a single or a few directions in the activations (Panickssery et al., 2023; Arditi et al., 2024) and weights (Wei et al., 2024; Jain et al., 2024), indicating that a small low-rank modification may be sufficient to induce safe behavior, without altering the entire weight space. Thus, we conjecture that the degradation in reasoning performance is caused by full-model finetuning introducing unnecessary changes in many directions, which interfere with critical weights responsible for reasoning.
To address this, we consider Low-Rank Adaptation (LoRA) (Hu et al., 2022), originally proposed as a parameter-efficient finetuning method to reduce training cost and memory usage. Rather than updating the full weight matrices, LoRA injects trainable low-rank matrices into existing layers while keeping the original weights frozen. Formally, a weight matrix ${\bm{W}}\in\mathbb{R}^{d\times k}$ is modified as:
@@
LoRA is particularly well-suited to our needs: it restricts updates to a low-rank subspace, thereby significantly reducing interference with the original weights. We will show in our experiments that this method works excellently, enabling the model to become safe while maintaining strong reasoning performance across benchmarks. As an additional benefit, LoRA is significantly more computationally efficient than full-model finetuning.
@@ 4 LoRA Bypasses the “Safety Tax”
In this section, we first introduce our safety finetuning and evaluation pipeline. Then, we evaluate models’ safety alignment and reasoning performance after full-model and LoRA safety finetuning.
@@ 4.1 Safety Alignment Fine-Tuning of Reasoning LLMs
We begin with a reasoning-capable language model and perform safety-alignment finetuning via SFT on a dataset of harmful questions paired with refusal responses, training the model to reject harmful requests. Following (Huang et al., 2025), we choose SFT over RL-based methods because it is simpler, more cost-effective, and avoids the need for additional components such as a reward model.
In our training setup, we compare (1) full-model finetuning, as in (Huang et al., 2025), where all model parameters are updated using standard gradient-based optimization; and (2) LoRA finetuning described in Section 3.
4.2 Evaluation of the Finetuned Model
After safety alignment finetuning is completed, we evaluate two aspects of the model: (1) safety, which is assessed using a dataset of harmful questions. We sample responses from the model for these questions and use Llama-Guard-3-8B—a model specialized in safety evaluation and shown to be the strongest safety evaluator in (Jiang et al., 2025)—to determine whether the responses are safe. The safety score is defined as the proportion of questions for which the model’s response is judged to be harmful. (2) reasoning ability, evaluated using multiple standard benchmark datasets containing questions on math, science, and coding—widely used to assess models’ reasoning capabilities. We consider the commonly used metric Pass@1 to measure accuracy on these benchmarks. For each question, we sample $n$ responses, compute the fraction of correct responses, and then average this accuracy over all questions. We set $n=8$.